# Phased Instruction Fine-Tuning for Large Language Models

### v1

@@ A.1 Histogram and density curve for Alpaca, Alpaca-cleaned and LIMA
A.2 ChatGPT-4 for evaluating win rate
@@
A.3 Experiment 4: Comparison win rate of IFT only on the Alpaca stage3 data vs. IFT on all the 52k Alpaca
Table 5: Win rate comparison of One-off IFT only on the third stage selected from Alpaca-3-stages against One-off IFT on the original Alpaca 52K data, using Llama-2 7B model. Each entry is the same as in Table 2.
Self-Instruct WizardLM Koala Vicuna OASST Anthropic
+0.79 +6.65 +2.50 +5.00 +3.81 +4.75
We compared the win rates of IFT conducted solely on the third stage, which includes 6.8K samples, against IFT performed on the entire 52K Alpaca data, as shown in Table 5. It was observed that the win rate for IFT on just stage 3 exceeds that of the full 52K dataset, aligning with conclusions drawn in Alpagasus Chen et al. (2024) that the quality of the instruction dataset is crucial. However, a comparison with Experiment 3 in Table 4 reveals that the win rate for IFT solely on stage 3 is significantly lower than when Phased IFT is conducted across all three stages, roughly only half of the latter’s. This underscores the existence of the progressive alignment hypothesis and validates the effectiveness of the Phased IFT approach.

### v2

@@ A.2 Histogram and density curve for Alpaca, Alpaca-cleaned and LIMA
A.3 ChatGPT-4 for evaluating win rate
@@
A.4 Comparison of win rate trends across multi-stages of uptraining
Figure 8 is a summary of this paper, which incorporates the results from Table 3. As can be seen from the figure, with the progression of uptraining, it demonstrates the winning rate growth trend of five LLMs on multi-stage sub-datasets with increasing difficulty. This forms a stark contrast to the winning rate trend of the same five LLMs on multi-stage sub-datasets with randomly distributed difficulty levels.
To be specific, the winning rate steadily continues to increase when uptraining on the sub-datasets with incremental difficulty. Conversely, when uptraining on the sub-datasets with random difficulty, the winning rate first increases then decreases, apart from the continuously declining winning rate trend of Llama3 70B. This indicates that the Progressive Alignment Hypothesis we proposed is an effective instruction learning method. It gradually learns from the pre-training model’s ability to predict the next word to eventually acquire the ability to complete the human end task.